# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v2

@@ 4.6 The GSMF and potential systematics
While it is exciting to obtain measurements down to $10^{5.4}{\rm M}_{\sun}$ and lower limits of $10^{3.8}{\rm M}_{\sun}$ for blue galaxies, caution is warranted, as extended objects in DECaLS are not guaranteed to be galaxies. For example, star-forming regions and spiral arms may be deblended into separate pieces. It is challenge to differentiate between multiple overlapping galaxies and a single structured galaxy (Beard et al., 1990; Dawson et al., 2016; Aihara et al., 2018b). If the density of fragmented sources is significant, it could lead to an overestimation of the GSMF at the low mass end, particularly for blue galaxies. To quantify and correct for these spurious sources, we mask regions in DECaLS within $2R_{e}$ of massive BGS galaxies and subsequently recalculate $\bar{n}_{2}w_{{\rm{p}}}$, where $R_{e}$ is the effective radius of the galaxy. Since $2R_{e}$ encompasses most of the galaxy’s light, this masking should effectively eliminate fake structures while minimizing impact on genuine neighboring galaxies. We test three different mass thresholds: $M_{*}^{\rm{mask}}=10^{10.0},10^{9.0},10^{8.0}\,{\rm M}_{\sun}$. Regions around BGS galaxies with $M_{*}>M_{*}^{\rm{mask}}$ are masked in DECaLS. Using $M_{*}^{\rm{mask}}=10^{8.0}\,{\rm M}_{\sun}$ results in the masking of regions around most BGS galaxies. We recalculate $\bar{n}_{2}w_{{\rm{p}}}$ only for ${\rm pop}_{2}$ with $M_{*}<M_{*}^{\rm{mask}}$, as the host galaxies are expected to be significantly more massive than any deblended substructures.
@@
From Figure 14, we observe that red galaxies dominate the high-mass end at $M_{*}>10^{10.6}{\rm M}_{\sun}$, while blue galaxies become increasingly prominent at lower masses, overtaking red galaxies at around $10^{9.8}{\rm M}_{\sun}$. However, the density of red galaxies rises sharply below $10^{8.6}{\rm M}_{\sun}$, once again dominating at $M_{*}<10^{7.6}{\rm M}_{\sun}$. This increasing trend persists down to the stellar mass limit of $10^{6.4}{\rm M}_{\sun}$ for red galaxies. If this trend holds beyond our observational limits, dwarf galaxies would predominantly be red. Peng et al. (2010) provided a physical explanation for the red GSMF, where galaxies with $M_{*}>M_{c}$ are primarily influenced by "mass quenching" in central galaxies, and those with $M_{*}<M_{c}$ are driven by "environmental quenching" in satellite galaxies. However, while their model predicted a slight upturn in the red GSMF around $10^{9.0}{\rm M}_{\sun}$, it is not nearly as steep as what we observe. This discrepancy suggests that the quenching mechanism for red dwarf galaxies might differ from our current understanding. Investigating this further is an interesting and important problem for future studies.
@@ 4.7 Comparison with previous nearby GSMFs
We also compare our results with those from GAMA DR4 at $z<0.1$ (Driver et al., 2022). To ensure a meaningful comparison, it is crucial to account for systematic errors in stellar mass, as we utilize different SED templates and photometry compared to GAMA. Therefore, we match GAMA galaxies used in (Driver et al., 2022) to DECaLS and recompute their stellar masses using our SED template and DECaLS photometry. A comparison between the two stellar mass estimates is shown in the left panel of Figure 16. We find that our stellar mass calculations are slightly lower than those from GAMA, with the difference increasing toward lower stellar masses. This trend follows a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=1.072\log_{10}(M^{\rm{GAMA}}_{*}/10^{10}{\rm M}_{\sun})-0.237$. The increasing discrepancy at the low mass end is expected, as most SED templates are calibrated at high masses. Recently, de los Reyes et al. (2024) highlighted that stellar mass estimates for low mass galaxies are sensitive to the assumed star formation history in the template, which can lead to differences of up to 0.4 dex. Despite this stellar mass discrepancy, we can still make a reasonable comparison after calibrating the GAMA stellar mass using the relation in Figure 16. We adjust the GSMF from (Driver et al., 2022) accordingly and present the comparison with our results in Figure 15. Additionally, since many studies utilize the Driver et al. (2022) stellar mass function, we also present our GSMFs calibrated to the Driver et al. (2022) stellar mass in Appendix B, allowing for easier comparison with other studies.
@@
We find that our results are in good agreement with Driver et al. (2022) for stellar masses in the range $10^{9.0}{\rm M}_{\sun}<M_{*}<10^{11.0}{\rm M}_{\sun}$. However, we observe slight discrepancies for $M_{*}>10^{11.0}{\rm M}_{\sun}$ and an increasing divergence below $10^{9.0}{\rm M}_{\sun}$. Both discrepancies are likely attributed to cosmic variance. At the high mass end, the limited survey area of GAMA, which covers only about 250 deg2, results in a scarcity of massive galaxies at $z<0.1$ available for calculating the GSMF, leading to significant uncertainties. In contrast, as shown as green dots in Figure 15, when we compare our GSMF with that from the much larger DESI Y1 BGS Bright sample at $z<0.2$ using the same SED template and $V_{\rm{max}}$ method (S. Moore 2025, in preparation), we find strong consistency up to $10^{11.8}{\rm M}_{\sun}$. This indicates that the discrepancies with GAMA for massive galaxies might be primarily due to its small sample size. Another possible factor could be differences in scatter in the estimated stellar masses, which could affect the GSMF at the steep, high-mass end. At the low mass end, the depth of GAMA is only $r_{\rm{KiDS}}\sim 19.65$, which is much shallower than DECaLS with $r\sim 23.0$. As a result, most low mass galaxies in GAMA are found in the local volume surrounding the Milky Way. It is well known that the Milky Way resides in a Local Void (LV) (Tully et al., 2008; Peebles & Nusser, 2010; Böhringer et al., 2020). Consequently, using the local volume to study the GSMF is likely to yield underestimated results. This discrepancy at $M_{*}<10^{9.0}M_{\odot}$ is also found when comparing to the BGS $V_{\rm{max}}$ results, which has the depth of $r=19.5$. Moreover, Driver et al. (2022) found that the GSMF from GAMA is broadly consistent with those from the Local Volume (Karachentsev & Makarova, 2019) and the Local Group (McConnachie, 2012) in the stellar mass range $10^{8.0}{\rm M}_{\sun}<M_{*}<10^{10.0}{\rm M}_{\sun}$. We reproduce their measurements in the Local Group in Figure 15 (brown triangles), applying the same stellar mass calibration in Figure 16 used for GAMA, under the assumption that Driver et al. (2022) made a consistent comparison.
@@
To quantify the bias from the LV, Chen et al. (2019) provide a comprehensive correction approach using the ELUCID simulation (Wang et al., 2014b, 2016b). This simulation was conducted using the reconstructed initial density field of the nearby universe within a box size of $500h^{-1}{\rm{Mpc}}$. They observed a rapid decrease in halo and galaxy densities toward lower redshifts, beginning from $z=0.03$. They derived the correction for the SDSS GSMF ($r<17.6$) by comparing the GSMF from the simulated SDSS region to those across the entire simulation box. Additionally, they examined corrections for samples with $r$-band magnitude limits ranging from $16$ to $20$. Recently, Wang et al. (2024) found that the corrected low mass end of the SDSS GSMF aligns well with results from their deeper $r=19.5$ sample, which combines DESI Y1 BGS and DESI Legacy Imaging Surveys samples. This indicates that the cosmic variance correction from Chen et al. (2019) is highly effective. Additionally, Wang et al. (2024) applied the correction for $r=19.5$ from Chen et al. (2019) to their GSMF, resulting in a much steeper increase at lower masses. We adopt the correction $f_{2}$ from Wang et al. (2024) and apply it to the BGS $V_{\rm{max}}$ results. The LV bias corrected BGS $V_{\rm{max}}$ GSMF is also shown in Figure 15. We find that it now aligns well with our GSMF down to $10^{7.6}{\rm M}_{\sun}$.We note that the correction from Chen et al. (2019) was performed within the SDSS footprint, which does not exactly match the DESI Y1 BGS sample. Additionally, as the GSMF from GAMA shows good agreement with the BGS $V_{\rm{max}}$ results, despite covering different sky regions, this suggests that the bias related to the LV may depend primarily on redshift rather than specific sky areas. This factor was not fully explored by Chen et al. (2019) and warrants further investigation. Moreover, Chen et al. (2019) did not provide corrections of the LV bias for lower masses due to the resolution limits of the ELUCID simulation. Thus, whether the discrepancies observed at lower masses can be fully attributed to cosmic variance requires further investigation.
@@
We acknowledge that GAMA has applied a LSS correction using the DFTOOLS method described in Obreschkow et al. (2018). However, this approach is based on a simplified assumption. Specifically, it uses more massive galaxies—which are observable over a larger volume—to estimate density variance. It then assumes that this variance applies equally to low-mass galaxies, adjusting their number densities accordingly. This assumption overlooks key complexities. It relies on the notion of linear galaxy bias, where galaxy overdensity is assumed to respond linearly to the matter overdensity. This is only valid when comparing galaxy samples in very large volumes with little mass variation $\delta_{g}\ll 1$. The local volume is not large enough and involves numerous non-linear processes. For instance, Figure 9 of Chen et al. (2019) shows, using a constrained simulation, the halo and galaxy number density change below $z<0.03$ differs significantly across mass bins. We would expect an even more complex behavior when considering how galaxy colour responds to environmental density. Furthermore, the DFTOOLS LSS correction requires normalization to determine the absolute correction factor. In GAMA’s case, this was done by fixing the total stellar mass within the survey volume at z<0.08 (Driver et al., 2022). However, since the region at z<0.03 is underdense, this approach likely underestimates the true correction.
@@ 4.8 Comparison with COSMOS GSMFs at higher redshifts
Since the LV correction from Chen et al. (2019) is highly model-dependent, we further validate our results by comparing our GSMF at $z<0.2$ with measurements from COSMOS-Web (Shuntov et al., 2025) at photometric redshifts $0.50.2<z_{\rm p}<$ 0.5 and $0.80.5<z_{\rm p}<$ 0.8. COSMOS-Web (Casey et al., 2023) combines deep ground- and space-based photometric data across 33 bands in the COSMOS field (Scoville et al., 2007), including recent observations from JWST (Gardner et al., 2006), and provides photometric redshifts and physical properties for galaxies over an area of approximately 0.431 deg2 (Shuntov et al., 2025). This region currently offers the only reliable constraints on the GSMF at $M_{*}<10^{9.0}{\rm M}_{\sun}$ beyond the local universe. To ensure a fair comparison, we calibrate the COSMOS-Web stellar masses to our scale. As the COSMOS-Web catalogue has not yet been published, we instead use the COSMOS2020 catalogue (Weaver et al., 2022), given the consistency of the COSMOS data reduction pipelines. We select all galaxies in COSMOS2020 with photometric redshifts $0.80.2<z_{\rm p}<$ 0.8 and compute their stellar masses using our GIGALE SED templates (Section 2.1), based solely on HSC $grz$ photometry for optimal consistency. These results are then compared to the COSMOS2020 stellar masses derived with Le Phare, as shown in the right panel of Figure 16. We find that our stellar masses are quite consistent with those from COSMOS2020, and their relation can be described by $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}\,M_{\odot})=0.948\,\log_{10}(M^{\rm{COSMOS}}_{*}/10^{10}\,M_{\odot})-0.012$. The comparison of GSMFs is presented in Figure 17. We find that our GSMF agrees well with the COSMOS-Web measurements down to $10^{7.6}{\rm M}_{\sun}$ at the two higher redshifts, within the uncertainties. This further indicate a potential higher number density of dwarf galaxies than reported in the local volume. Shuntov et al. (2025) did not provide separate GSMFs for blue and red galaxies, but the sharp rise in the red GSMF at $M_{*}<10^{8.5}{\rm M}_{\sun}$ is evident ($\alpha_{2}\sim-2$) in the COSMOS2020 results (Weaver et al., 2023) for $z<0.8$. We plan to explore this further once the full COSMOS-Web catalogues are released.
In Figure 17, we also reproduce the GSMF from G10-COSMOS by Wright et al. (2017) (see also Driver et al. 2018; Wright et al. 2018), based on a combination of spectroscopic and photometric redshifts at $z_{\rm p}<0.1$ (green triangles). The stellar masses are calibrated using the relation for GAMA, under the assumption that Wright et al. (2017) applied a consistent methodology to G10-COSMOS as was used for GAMA, as detailed in Driver et al. (2018). The results from G10-COSMOS appear to lie between those from GAMA and COSMOS-Web/our work. However, caution is warranted when using photometric redshifts at such low redshifts.

### v3

@@ 4.6 The GSMF and potential systematics
While it is exciting to obtain measurements down to $10^{5.4}{\rm M}_{\sun}$ and lower limits of $10^{3.8}{\rm M}_{\sun}$ for blue galaxies, caution is warranted, as extended objects in DECaLS are not guaranteed to be galaxies. For example, star-forming regions and spiral arms may be deblended into separate pieces. It is challenge to differentiate between multiple overlapping galaxies and a single structured galaxy (1990MNRAS.247..311B; 2016ApJ...816...11D; 2018PASJ...70S...8A). If the density of fragmented sources is significant, it could lead to an overestimation of the GSMF at the low mass end, particularly for blue galaxies. To quantify and correct for these spurious sources, we mask regions in DECaLS within $2R_{e}$ of massive BGS galaxies and subsequently recalculate $\bar{n}_{2}w_{{\rm{p}}}$, where $R_{e}$ is the effective radius of the galaxy. Since $2R_{e}$ encompasses most of the galaxy’s light, this masking should effectively eliminate fake structures while minimizing impact on genuine neighboring galaxies. We test three different mass thresholds: $M_{*}^{\rm{mask}}=10^{10.0},10^{9.0},10^{8.0}\,{\rm M}_{\sun}$. Regions around BGS galaxies with $M_{*}>M_{*}^{\rm{mask}}$ are masked in DECaLS. Using $M_{*}^{\rm{mask}}=10^{8.0}\,{\rm M}_{\sun}$ results in the masking of regions around most BGS galaxies. We recalculate $\bar{n}_{2}w_{{\rm{p}}}$ only for ${\rm pop}_{2}$ with $M_{*}<M_{*}^{\rm{mask}}$, as the host galaxies are expected to be significantly more massive than any deblended substructures.
@@
From Figure 14, we observe that red galaxies dominate the high-mass end at $M_{*}>10^{10.6}{\rm M}_{\sun}$, while blue galaxies become increasingly prominent at lower masses, overtaking red galaxies at around $10^{9.8}{\rm M}_{\sun}$. However, the density of red galaxies rises sharply below $10^{8.6}{\rm M}_{\sun}$, once again dominating at $M_{*}<10^{7.6}{\rm M}_{\sun}$. This increasing trend persists down to the stellar mass limit of $10^{6.4}{\rm M}_{\sun}$ for red galaxies. If this trend holds beyond our observational limits, dwarf galaxies would predominantly be red. 2010ApJ...721..193P provided a physical explanation for the red GSMF, where galaxies with $M_{*}>M_{c}$ are primarily influenced by "mass quenching" in central galaxies, and those with $M_{*}<M_{c}$ are driven by "environmental quenching" in satellite galaxies. However, while their model predicted a slight upturn in the red GSMF around $10^{9.0}{\rm M}_{\sun}$, it is not nearly as steep as what we observe. This discrepancy suggests that the quenching mechanism for red dwarf galaxies might differ from our current understanding. Investigating this further is an interesting and important problem for future studies.
@@ 4.7 Comparison with previous nearby GSMFs
We also compare our results with those from GAMA DR4 at $z<0.1$ (2022MNRAS.513..439D). To ensure a meaningful comparison, it is crucial to account for systematic errors in stellar mass, as we utilize different SED templates and photometry compared to GAMA. Therefore, we match GAMA galaxies used in (2022MNRAS.513..439D) to DECaLS and recompute their stellar masses using our SED template and DECaLS photometry. A comparison between the two stellar mass estimates is shown in the left panel of Figure 16. We find that our stellar mass calculations are slightly lower than those from GAMA, with the difference increasing toward lower stellar masses. This trend follows a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=1.072\log_{10}(M^{\rm{GAMA}}_{*}/10^{10}{\rm M}_{\sun})-0.237$. The increasing discrepancy at the low mass end is expected, as most SED templates are calibrated at high masses. Recently, 2024arXiv240903959D highlighted that stellar mass estimates for low mass galaxies are sensitive to the assumed star formation history in the template, which can lead to differences of up to 0.4 dex. Despite this stellar mass discrepancy, we can still make a reasonable comparison after calibrating the GAMA stellar mass using the relation in Figure 16. We adjust the GSMF from (2022MNRAS.513..439D) accordingly and present the comparison with our results in Figure 15. Additionally, since many studies utilize the 2022MNRAS.513..439D stellar mass function, we also present our GSMFs calibrated to the 2022MNRAS.513..439D stellar mass in Appendix B, allowing for easier comparison with other studies.
@@
We find that our results are in good agreement with 2022MNRAS.513..439D for stellar masses in the range $10^{9.0}{\rm M}_{\sun}<M_{*}<10^{11.0}{\rm M}_{\sun}$. However, we observe slight discrepancies for $M_{*}>10^{11.0}{\rm M}_{\sun}$ and an increasing divergence below $10^{9.0}{\rm M}_{\sun}$. Both discrepancies are likely attributed to cosmic variance. At the high mass end, the limited survey area of GAMA, which covers only about 250 deg2, results in a scarcity of massive galaxies at $z<0.1$ available for calculating the GSMF, leading to significant uncertainties. In contrast, as shown as green dots in Figure 15, when we compare our GSMF with that from the much larger DESI Y1 BGS Bright sample at $z<0.2$ using the same SED template and $V_{\rm{max}}$ method (S. Moore 2025, in preparation), we find strong consistency up to $10^{11.8}{\rm M}_{\sun}$. This indicates that the discrepancies with GAMA for massive galaxies might be primarily due to its small sample size. Another possible factor could be differences in scatter in the estimated stellar masses, which could affect the GSMF at the steep, high-mass end. At the low mass end, the depth of GAMA is only $r_{\rm{KiDS}}\sim 19.65$, which is much shallower than DECaLS with $r\sim 23.0$. As a result, most low mass galaxies in GAMA are found in the local volume surrounding the Milky Way. It is well known that the Milky Way resides in a Local Void (LV) (2008ApJ...676..184T; 2010Natur.465..565P; 2020A&A...633A..19B). Consequently, using the local volume to study the GSMF is likely to yield underestimated results. This discrepancy at $M_{*}<10^{9.0}M_{\odot}$ is also found when comparing to the BGS $V_{\rm{max}}$ results, which has the depth of $r=19.5$. Moreover, 2022MNRAS.513..439D found that the GSMF from GAMA is broadly consistent with those from the Local Volume (2019Ap.....62..293K) and the Local Group (2012AJ....144....4M) in the stellar mass range $10^{8.0}{\rm M}_{\sun}<M_{*}<10^{10.0}{\rm M}_{\sun}$. We reproduce their measurements in the Local Group in Figure 15 (brown triangles), applying the same stellar mass calibration in Figure 16 used for GAMA, under the assumption that 2022MNRAS.513..439D made a consistent comparison.
@@
To quantify the bias from the LV, 2019ApJ...872..180C provide a comprehensive correction approach using the ELUCID simulation (2014ApJ...794...94W; 2016ApJ...831..164W). This simulation was conducted using the reconstructed initial density field of the nearby universe within a box size of $500h^{-1}{\rm{Mpc}}$. They observed a rapid decrease in halo and galaxy densities toward lower redshifts, beginning from $z=0.03$. They derived the correction for the SDSS GSMF ($r<17.6$) by comparing the GSMF from the simulated SDSS region to those across the entire simulation box. Additionally, they examined corrections for samples with $r$-band magnitude limits ranging from $16$ to $20$. Recently, 2024ApJ...971..119W found that the corrected low mass end of the SDSS GSMF aligns well with results from their deeper $r=19.5$ sample, which combines DESI Y1 BGS and DESI Legacy Imaging Surveys samples. This indicates that the cosmic variance correction from 2019ApJ...872..180C is highly effective. Additionally, 2024ApJ...971..119W applied the correction for $r=19.5$ from 2019ApJ...872..180C to their GSMF, resulting in a much steeper increase at lower masses. We adopt the correction $f_{2}$ from 2024ApJ...971..119W and apply it to the BGS $V_{\rm{max}}$ results. The LV bias corrected BGS $V_{\rm{max}}$ GSMF is also shown in Figure 15. We find that it now aligns well with our GSMF down to $10^{7.6}{\rm M}_{\sun}$.We note that the correction from 2019ApJ...872..180C was performed within the SDSS footprint, which does not exactly match the DESI Y1 BGS sample. Additionally, as the GSMF from GAMA shows good agreement with the BGS $V_{\rm{max}}$ results, despite covering different sky regions, this suggests that the bias related to the LV may depend primarily on redshift rather than specific sky areas. This factor was not fully explored by 2019ApJ...872..180C and warrants further investigation. Moreover, 2019ApJ...872..180C did not provide corrections of the LV bias for lower masses due to the resolution limits of the ELUCID simulation. Thus, whether the discrepancies observed at lower masses can be fully attributed to cosmic variance requires further investigation.
@@
We acknowledge that GAMA has applied a LSS correction using the DFTOOLS method described in 2018MNRAS.474.5500O. However, this approach is based on a simplified assumption. Specifically, it uses more massive galaxies—which are observable over a larger volume—to estimate density variance. It then assumes that this variance applies equally to low-mass galaxies, adjusting their number densities accordingly. This assumption overlooks key complexities. It relies on the notion of linear galaxy bias, where galaxy overdensity is assumed to respond linearly to the matter overdensity. This is only valid when comparing galaxy samples in very large volumes with little mass variation $\delta_{g}\ll 1$. The local volume is not large enough and involves numerous non-linear processes. For instance, Figure 9 of 2019ApJ...872..180C shows, using a constrained simulation, the halo and galaxy number density change below $z<0.03$ differs significantly across mass bins. We would expect an even more complex behavior when considering how galaxy colour responds to environmental density. Furthermore, the DFTOOLS LSS correction requires normalization to determine the absolute correction factor. In GAMA’s case, this was done by fixing the total stellar mass within the survey volume at z<0.08 (2022MNRAS.513..439D). However, since the region at z<0.03 is underdense, this approach likely underestimates the true correction.
@@ 4.8 Comparison with COSMOS GSMFs at higher redshifts
Since the LV correction from 2019ApJ...872..180C is highly model-dependent, we further validate our results by comparing our GSMF at $z<0.2$ with measurements from COSMOS-Web (2025A&A...695A..20S) at photometric redshifts $0.2<z_{\rm p}<0.5$ and $0.5<z_{\rm p}<0.8$. COSMOS-Web (2023ApJ...954...31C) combines deep ground- and space-based photometric data across 33 bands in the COSMOS field (2007ApJS..172...38S), including recent observations from JWST (2006SSRv..123..485G), and provides photometric redshifts and physical properties for galaxies over an area of approximately 0.431 deg2 (2025A&A...695A..20S). This region currently offers the only reliable constraints on the GSMF at $M_{*}<10^{9.0}{\rm M}_{\sun}$ beyond the local universe. To ensure a fair comparison, we calibrate the COSMOS-Web stellar masses to our scale. As the COSMOS-Web catalogue has not yet been published, we instead use the COSMOS2020 catalogue (2022ApJS..258...11W), given the consistency of the COSMOS data reduction pipelines. We select all galaxies in COSMOS2020 with photometric redshifts $0.2<z_{\rm p}<0.8$ and compute their stellar masses using our GIGALE SED templates (Section 2.1), based solely on HSC grzgrz photometry for optimal consistency. These results are then compared to the COSMOS2020 stellar masses derived with Le Phare, as shown in the right panel of Figure 16. We find that our stellar masses are quite consistent with those from COSMOS2020, and their relation can be described by $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}\,M_{\odot})=0.948\,\log_{10}(M^{\rm{COSMOS}}_{*}/10^{10}\,M_{\odot})-0.012$. The comparison of GSMFs is presented in Figure 17. We find that our GSMF agrees well with the COSMOS-Web measurements down to $10^{7.6}{\rm M}_{\sun}$ at the two higher redshifts, within the uncertainties. This further indicate a potential higher number density of dwarf galaxies than reported in the local volume. 2025A&A...695A..20S did not provide separate GSMFs for blue and red galaxies, but the sharp rise in the red GSMF at $M_{*}<10^{8.5}{\rm M}_{\sun}$ is evident ($\alpha_{2}\sim-2$) in the COSMOS2020 results (2023A&A...677A.184W) for $z<0.8$. We plan to explore this further once the full COSMOS-Web catalogues are released.
In Figure 17, we also reproduce the GSMF from G10-COSMOS by 2017MNRAS.470..283W (see also 2018MNRAS.475.2891D; 2018MNRAS.480.3491W), based on a combination of spectroscopic and photometric redshifts at $z_{\rm p}<0.1$ (green triangles). The stellar masses are calibrated using the relation for GAMA, under the assumption that 2017MNRAS.470..283W applied a consistent methodology to G10-COSMOS as was used for GAMA, as detailed in 2018MNRAS.475.2891D. The results from G10-COSMOS appear to lie between those from GAMA and COSMOS-Web/our work. However, caution is warranted when using photometric redshifts at such low redshifts.